OpenStudy (anonymous):

Three baseball fans estimated the top speed of a baseball pitcher’s fastball. The table shows their estimates. The top speed of the pitcher’s fastball is known to be 102 mph. Which statements are true? Name Speed guessed (mph) Lisa 97 Justin 93 Mohammed 105 Choose exactly two answers that are correct. A. Justin’s estimate had a lower percent error than Lisa’s. B. Mohammed’s estimate had the lowest percent error. C. Lisa’s estimate had about a 4.9% error. D. Justin’s estimate had a –9% percent error.

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

\[\frac{ 102-97 }{ 97 }\times100=?\]\[\frac{ 102-93 }{ 93 }\times100\]\[\frac{ 105-102 }{ 102 }\times100=?\] solve it

OpenStudy (anonymous):

@urbanmorgans ^

OpenStudy (anonymous):

here

OpenStudy (anonymous):

okay solve them :)

OpenStudy (anonymous):

1. 5/97 2.3/31 3.1/34

OpenStudy (anonymous):

@shifuyanli